# Quenched coalescent for diploid population models with selfing and overlapping generations

### v1

@@ 7.2 Convergence of the discrete-time ancestral graph
We begin by giving a construction of the discrete-time ancestral graph $G^{N,n}=\left(l^{N},m^{N},\xi^{N}\right)$ as a discrete-time process taking values in $\Theta$.
@@
That is, $\tilde{G}^{N,n}$ consists of a sequence of edges $\left(\tilde{E}_{k}^{N}\right)$ where the edges from $\tilde{V}_{k}^{N}$ to $\tilde{V}_{k+1}^{N}$ consist of those edges from individual to itself for persisting in that time-step (i.e. not dying) and also edges from parents to children for reproductive events in this time-step. The edge set $\tilde{E}_{k}^{N}$ is merely the edges of the pedigree connecting those elements of $\tilde{V}_{k}^{N}$.
@@
  • $m^{N}(k)$ is the index of the node in $\tilde{G}$ the subject of most recent outcrossing, i.e.
@@
    Note that while it is possible in our model that more than one outcrossing occurs in each time-step, this will occur with negligible probability over the lifespan of the graph. This is shown in Lemma 7.6.
@@
We now recall the definition of the haploid map $F_{\rm hap}$ of an element of $\mathcal{S}_{n}$, defined in (3.10). The haploid map $F_{\rm hap}$ maps $\mathcal{S}_{n}$ into $\mathcal{E}_{n}$ by collapsing all elements of the form $\{A,B\}$ into $A\cup B$. Formally, for
$\xi=\{(C_{1},C_{2}),\ldots,(C_{2x-1},C_{2x}),C_{2x+1},\ldots,C_{b}\}$
we define
$F_{\rm hap}\left(\xi\right):=\{C_{1}\cup C_{2},\ldots,C_{2x-1}\cup C_{2x},C_{2x+1},\ldots,C_{b}\}.$
@@
$Q^{N}_{n}=\left(q_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}=c_{N}^{-1}\left(\mathit{p}_{\xi\eta}^{N,n}-\delta_{\xi\eta}\right)_{\xi,\eta\in\mathcal{E}_{n}},$ (7.2)
where $\delta_{\xi\eta}$ is the Kronecker delta function. Recall too that the infinite population limit generator $Q_{n}=\left(q_{\xi\eta}^{n}\right)_{\xi,\eta\in\mathcal{E}_{n}}$. We calculate the one-step transition probabilities for $G^{N,n}$ explicitly in terms of $Q^{N}_{n}$ in the following lemma.
@@
Suppose that Assumption 3.4 holds. Then the one-step transition probabilities of the Markov chain $\{G^{N,n}(k)\}_{k\in\mathbb{Z}_{+}}$ satisfy the following asymptotic property: given that $G^{N,n}(k)=(l,m,\xi)\in\Theta$, it holds that
@@
for each $(m^{\prime},l^{\prime},l^{\prime\prime},\eta)$ such that $m^{\prime}\neq m$ and $l^{\prime}\neq l+1$, $l^{\prime\prime}\neq|\eta|$, and $\eta\in\mathcal{E}_{l}$. In the above, the error terms $O(c_{N}),o(c_{N}),O(d_{N}c_{N})$ are all uniform for $(l,m,\xi)\in\Theta$.
@@
Observe that the probability of either an outcrossing or a coalescence of the $l$ nodes is bounded above by $2c_{N}\binom{l}{2}+ld_{N}$. By Assumption 3.4, the probability of neither a coalescence nor an outcrossing is $1-l^{2}O(c_{N})$. This gives the first line.
Now we show that the probability that two outcrossings occur in the same time-step occurs with negligible probability, implying line 3. Indeed, the probability that two outcrossings occur is bounded by the probability that any of the $\binom{l}{2}$ possible pairs of particles are both children and both are the result of outcrossing. The probability that any such pair consists of two children is $\mathbb{E}\left[\frac{K_{N}(K_{N}-1)}{N(N-1)}\right]$. The probability that they both are the result of outcrossing is $(1-\alpha_{N})^{2}$. Therefore the probability of at least two outcrossings is at most
@@
This, together with the fact that $\alpha_{N}\to 1$ (by Assumption 3.4), implies that the probability of at least two outcrossings is bounded by $\binom{l}{2}o(d_{N})$. This gives the third line.
@@
which gives line $5$.
We now show that an outcrossing occurs during a coalescence with negligible probability, implying the fourth line. As $\xi_{0}^{l}$ consists of particles in $l$ distinct individuals, the coalescence rate is independent to the selfing rate. In particular, the probability that we have a coalescence with structure $\eta$ and an outcrossing is therefore $O((1-\alpha_{N})q^{N,l}_{\xi\eta})\subset o(c_{N})$. This gives the fourth line. ∎
@@
Suppose that, as $N\to\infty$, Assumptions 3.3, 3.4, and 3.6 hold. Then the sped-up discrete-time ancestral graph $\bar{G}^{N}=\left(G^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)\right)_{t\in\mathbb{R}_{+}}$ converges weakly in the $J_{1}$ topology in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to a $Q$-$\lambda$ ancestral graph $G^{n}_{Q,\lambda}$ with respect to the measures $\mathbb{P}_{\xi_{0}^{n}}^{N}$, where $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is as in Assumption 3.6.
@@
We proceed as in the proof of NWF (25)[Lemma 7.6].
Let $T^{(N)}$ denote the linear operator on the space $C_{b}\left(\Theta\right)$ of bounded continuous functions on $\Theta$ defined by $T^{(N)}f(l,m,\xi)=\mathbb{E}\left[f\left(G^{N}(1)\right)\mid G^{N,n}(0)=(l,m,\xi)\right]$. The generator $\mathscr{L}^{N}$ of the discrete-time process $G^{N,n}$ is given by
$\mathscr{L}^{N}f(l,m,\xi)=\left(T^{(N)}-I\right)f(l,m,\xi),$
which can be calculated to sufficient accuracy as in Equation 7.3. Let $\mathscr{L}$ be the infinitesimal generator of the continuous-time process $G_{Q,\lambda}^{n}:=\left(G_{Q,\lambda}^{n}(t)\right)_{t\in\mathbb{R}_{+}}$. That is, by Equation (7.1),
$\mathscr{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}\tilde{q}_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.4)
It follows from Equation (7.3) that for all $f\in C_{b}(\Theta)$ with finite support that,
$\sup_{(l,m,\xi)\in\Theta}\left|c_{N}^{-1}\mathscr{L}^{N}f(l,m,\xi)-\mathscr{L}f(l,m,\xi)\right|\to 0\text{ as $N\to\infty$.}$
Let $\{T(t)\}_{t\in\mathbb{R}_{+}}$ be the semigroup on $C_{b}(\Theta)$ of $G_{Q,\lambda}^{n}$. By EK (09)[Theorem 6.6, Chapter 1] and (7.4), it holds that
$\lim_{N\to\infty}\sup_{0\leq t\leq T}\sup_{(l,m,\xi)\in\Theta}\left|\left(T^{(N)}\right)^{\lfloor tc_{N}^{-1}\rfloor}f(l,m,\xi)-T(t)f(l,m,\xi)\right|=0$ (7.5)
for all $f$ in the domain of $\mathscr{L}$
We now proceed to demonstrate tightness. Observe that, for any $n$ that the set $K_{n}:=[n]\times[n]\times\cup_{2\leq k\leq n}\mathcal{E}_{k}$ is compact. To establish the compact containment condition EK (09)[(7.9), p.129] holds if for any $\varepsilon\in(0,1)$ and $T\in(0,\infty)$, there exists $M(\varepsilon,T)\in\mathbb{Z}_{+}$ such that
@@
Notice that for all time $t$ we have that
$|\xi^{N}(t)|,|m^{N}(t)|\leq\sup_{0\leq s\leq t}l^{N}(t).$
@@
see KL (98)[Lemma 5.1] or EK (09)[Proposition 4.1.7]. A truncation argument will enable us to take $f$ to be the identity function $Id$ (i.e. when $Id(n)=n$ for all $n\in\mathbb{Z}_{+}$) to obtain
@@
Finally, (7.7) can be obtained by the same argument above, when $L_{t}$ and $\int_{0}^{t}\mathcal{Q}f(L_{s})ds$ are replaced, respectively, by $l^{N}(k)$ and $\sum_{i=0}^{k-1}\mathcal{Q}^{N}f(l^{N}(i))$ for $\mathcal{Q}^{N}$ the generator of $l^{N}$. Indeed, the uniform control in $l$ given by Lemma 7.6, which depends on Assumptions 3.3 and 3.4, is enough for the argument above to follow as Equation 7.11 holds with the limiting transition rates $q^{n}_{\xi\eta}$ replaced by the finite $N$ transition rates $q^{N,n}_{\xi\eta}$, which follows by Assumption 3.6 (here $\xi,\eta$ are simply placeholders for the true elements used in the above equation). (7.7) implies the convergence of $\bar{G}^{N}$ in $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$ to $G_{\tilde{Q},\lambda}^{n}$ by EK (09)[Corollary 8.9, Chapter 4] and (7.5).
The discrete-time ancestral graph $G^{N,n}=(G^{N,n}(k))_{k\in\mathbb{Z}_{+}}$ can be viewed graphically as having $n$ initial nodes (the nodes at the bottom of Figure 6). These $n$ nodes correspond to the $n$ individuals $\{\hat{X}_{i}^{N}(0)\}_{1\leq i\leq n}$ from whom we sampled the $n$ lineages at time-step 0. Viewed as a $\Theta$-valued process, this graph has the following key advantages: it get rid of the labels of all individuals, while keeping track of the set of all possible ancestral individuals (of the sample) and their parental relationships.
@@
Define $x_{i}^{N}(0):=i$ for all $1\leq i\leq n$. Suppose $x_{i}^{N}(j)$ is defined for all $j\leq k$ for induction. If $\hat{X}_{i}^{N}(k)$ experiences an outcrossing in the $k$th time-step, then $x_{i}^{N}(k+1)$ is equal to $l^{N}(k+1)$ with probability $\frac{1}{2}$ or else remains equal to $x_{i}^{N}(k)$ with probability $\frac{1}{2}$. (In the event that there is more than one outcrossing in a single time-step, it may be that each outcrossing individual coalesces in the position $l+1$; in the limited outcrossing regime, this possibility is negligible, so we do not need to consider this case in any detail.) We therefore have $x_{i}^{N}=\left(x_{i}^{N}(k)\right)_{k\in\mathbb{Z}_{+}}$ defined for all time. If any two walks $x_{i}^{N}$ and $x_{j}^{N}$ coalesce at any point in time, they are then required to remain identical for all time thereafter. This gives a family of $n$ simple random walks on $G^{N,n}$ satisfying
@@
Lemma 7.8 is readily extended to the joint convergence of the graph with the walks thereon, namely the following lemma.
@@
We may also define an associated coalescent process based on these $x_{i}^{N}$. Define $\mathcal{L}^{N,n}_{G}$ to be the quenched law of the sped-up coalescent process $\Pi_{G}^{N,n}=\left(\Pi_{G}^{N,n}(t)\right)_{t\in\mathbb{R}_{+}}$ defined by
$i\sim j\text{ with respect to }\Pi_{G}^{N,n}(t)\text{ if }x_{i}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)=x_{j}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right).$
That is
@@
Note that $\mathcal{L}^{N,n}_{G}$ defined in (7.18) is equal to $\mathbb{P}\left(\chi^{N,n}(\lfloor tc_{N}^{-1}\rfloor)\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ which admits no random time change, unlike that of $\mathcal{L}^{N,n}$. This is because, unlike the random walks on the pedigree, we assume being in the same node in the ancestral graph $G^{N,n}$ means instantaneous coalescence of the particles. The comparison between $\mathcal{L}^{N,n}_{G}$ and $\mathcal{L}^{N,n}$ which squares our convergence results with this discrepancy is Lemma 7.14.
We will show in Lemma 7.14 that $\mathcal{L}^{N,n}$ and $\mathcal{L}^{N,n}_{G}$ have the same weak limit. A suitable notion of continuity for random walks on $\Theta$-valued processes, established in Section 7.3, will then give the claimed result.
7.3 Random walks on $\Theta$-valued processes
To each element $g=\left(g(t)\right)_{t\in\mathbb{R}_{+}}=\left(l(t),m(t),\xi(t)\right)_{t\in\mathbb{R}_{+}}$ in the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},\Theta\right)$, we will define an associated coalescent process $\chi(g)$. Let $g(t-)=\lim_{s\uparrow t}g(s)=(l(t-),m(t-),\xi(t-))$ denote the left limit of $g$ at $t$. We suppose that $g(0)=(l_{0},m_{0},\xi_{0})$ and define, for all $1\leq i\leq l_{0}$ that $x_{i}(0)=i$. The $x_{i}$ may be extended to all time as follows:

### v2

@@ 7.2 Convergence of the discrete-time ancestral graph
We begin by giving a construction of the discrete-time ancestral graph $G^{N,n}=\left(l^{N},m^{N},\xi^{N}\right)$ as a discrete-time process taking values in $\Omega$.
@@
That is, $\tilde{G}^{N,n}$ consists of a sequence of edges $\left(\tilde{E}_{k}^{N}\right)$ where the edges from $\tilde{V}_{k}^{N}$ to $\tilde{V}_{k+1}^{N}$ consist of those edges from individual to itself for persisting in that time-step (i.e. not dying) and also edges from parents to offspring for reproductive events in this time-step. The edge set $\tilde{E}_{k}^{N}$ is merely the edges of the pedigree connecting those elements of $\tilde{V}_{k}^{N}$.
@@
  • $m^{N}(k)$ is the index of the node in $\tilde{G}$ the subject of most recent outcrossing event, i.e.
@@
    Note that while it is possible in our model that more than one outcrossing event occurs in each time-step, this will occur with negligible probability over the lifespan of the graph. This is shown in Lemma 7.6.
@@
$Q^{N}_{n}=\left(q_{\xi\eta}^{N,n}\right)_{\xi,\eta\in\mathcal{E}_{n}}=c_{N}^{-1}\left(h_{\xi\eta}^{N,n}-\delta_{\xi\eta}\right)_{\xi,\eta\in\mathcal{E}_{n}},$ (7.3)
where $\delta_{\xi\eta}$ is the Kronecker delta function. We calculate the one-step transition probabilities for $G^{N,n}$ explicitly in terms of $Q^{N}_{n}$ in the following lemma.
@@
Suppose that Assumption 3.5 holds. Then the one-step transition probabilities of the Markov chain $\{G^{N,n}(k)\}_{k\in\mathbb{Z}_{+}}$ satisfy the following asymptotic property: given that $G^{N,n}(k)=(l,m,\xi)\in\Omega$, it holds that
@@
for each $(m^{\prime},l^{\prime},l^{\prime\prime},\eta)$ such that $m^{\prime}\neq m$ and $l^{\prime}\neq l+1$, $l^{\prime\prime}\neq|\eta|$, and $\eta\in\mathcal{E}_{l}$. In the above, the error terms $O(c_{N}),o(c_{N}),O(d_{N}c_{N})$ are all uniform for $(l,m,\xi)\in\Omega$.
@@
Observe that the probability of either an outcrossing event or a coalescent event among the $l$ nodes is bounded above by $2c_{N}\binom{l}{2}+ld_{N}$. By Assumption 3.5, the probability of neither an outcrossing event nor a coalescent event is $1-l^{2}O(c_{N})$. This gives the first line.
Now we show that the probability that two outcrossing events occur in the same time-step occurs with negligible probability, implying line 3. Indeed, the probability that two outcrossing events occur is bounded by the probability that any of the $\binom{l}{2}$ possible pairs of particles are both offspring and both are the result of outcrossing. The probability that any such pair consists of two offspring is $\mathbb{E}\left[\frac{K_{N}(K_{N}-1)}{N(N-1)}\right]$. The probability that they both are the result of outcrossing is $(1-\alpha_{N})^{2}$. Therefore the probability of at least two outcrossing events is at most
@@
This, together with the fact that $\alpha_{N}\to 1$ (by Assumption 3.5), implies that the probability of at least two outcrossing events is bounded by $\binom{l}{2}o(d_{N})$. This gives the third line of (7.4).
@@
which gives the fifth line of (7.4).
We now show that an outcrossing event occurs during a coalescence with negligible probability, implying the fourth line. As $\xi_{0}^{l}$ consists of particles in $l$ distinct individuals, the coalescence rate is independent to the selfing rate. In particular, the probability that we have a coalescence with structure $\eta$ and an outcrossing event is therefore $O((1-\alpha_{N})q^{N,l}_{\xi\eta})\subset o(c_{N})$. This gives the fourth line of (7.4). ∎
@@
Suppose that Assumptions 3.4, 3.5, and 3.7 hold. Then the sped-up discrete-time ancestral graph $\bar{G}^{N}=\left(G^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)\right)_{t\in\mathbb{R}_{+}}$ converges weakly in the $J_{1}$ topology in $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$ to a $Q$-$\lambda$ ancestral graph $G^{n}_{Q,\lambda}$ with respect to the measures $\mathbb{P}_{\xi_{0}^{n}}^{N}$, where $Q=\left(Q_{n}\right)_{n\in\mathbb{N}}$ is as in Assumption 3.7.
@@
We proceed as in the proof of (NWF, 25, Lemma 7.6). Let $T^{(N)}$ denote the linear operator on the space $C_{b}\left(\Omega\right)$ of bounded continuous functions on $\Omega$ defined by $T^{(N)}f(l,m,\xi)=\mathbb{E}\left[f\left(G^{N}(1)\right)\mid G^{N,n}(0)=(l,m,\xi)\right]$. The generator $\mathfrak{L}^{N}$ of the discrete-time process $G^{N,n}$ is given by
$\mathfrak{L}^{N}f(l,m,\xi)=\left(T^{(N)}-I\right)f(l,m,\xi),$
which can be calculated to sufficient accuracy as in Equation 7.4. Let $\mathfrak{L}$ be the infinitesimal generator of the continuous-time process $G_{Q,\lambda}^{n}:=\left(G_{Q,\lambda}^{n}(t)\right)_{t\in\mathbb{R}_{+}}$. That is, by Equation (7.2),
$\mathfrak{L}f(l,m,\xi)=\sum_{j=1}^{l}\lambda\left[f(l+1,j,\xi)-f(l,m,\xi)\right]+\sum_{\eta\in\mathcal{E}_{l}\setminus\{\xi_{0}^{l}\}}\tilde{q}_{\xi_{0}^{l}\eta}^{l}\left[f(|\eta|,m,\eta)-f(l,m,\xi)\right]$ (7.5)
It follows from Equation (7.4) that for all $f\in C_{b}(\Omega)$ with finite support that,
$\sup_{(l,m,\xi)\in\Omega}\left|c_{N}^{-1}\mathfrak{L}^{N}f(l,m,\xi)-\mathfrak{L}f(l,m,\xi)\right|\to 0\text{ as $N\to\infty$.}$
Let $\{T(t)\}_{t\in\mathbb{R}_{+}}$ be the semigroup on $C_{b}(\Omega)$ of $G_{Q,\lambda}^{n}$. By (EK, 09, Theorem 6.6, Chapter 1) and (7.5), it holds that
$\lim_{N\to\infty}\sup_{0\leq t\leq T}\sup_{(l,m,\xi)\in\Omega}\left|\left(T^{(N)}\right)^{\lfloor tc_{N}^{-1}\rfloor}f(l,m,\xi)-T(t)f(l,m,\xi)\right|=0$ (7.6)
for all $f$ in the domain of $\mathfrak{L}$
We now proceed to demonstrate tightness. Observe that, for any $n$ that the set $K_{n}:=[n]\times[n]\times\cup_{2\leq k\leq n}\mathcal{E}_{k}$ is compact. To establish the compact containment condition (EK, 09, (7.9), p.129) holds if for any $\varepsilon\in(0,1)$ and $T\in(0,\infty)$, there exists $M(\varepsilon,T)\in\mathbb{Z}_{+}$ such that
@@
Notice that for all time $t\in\mathbb{R}_{+}$,
$\max\left\{|\xi^{N}(t)|,\,|m^{N}(t)|\right\}\leq\sup_{0\leq s\leq t}l^{N}(t).$
@@
see (KL, 98, Lemma 5.1) or (EK, 09, Proposition 4.1.7). A truncation argument will enable us to take $f$ to be the identity function $Id$ (i.e. when $Id(n)=n$ for all $n\in\mathbb{Z}_{+}$) to obtain
@@
Finally, (7.8) can be obtained by the same argument above, when $L_{t}$ and $\int_{0}^{t}\mathcal{Q}f(L_{s})ds$ are replaced, respectively, by $l^{N}(k)$ and $\sum_{i=0}^{k-1}\mathcal{Q}^{N}f(l^{N}(i))$ for $\mathcal{Q}^{N}$ the generator of $l^{N}$. Indeed, the uniform control in $l$ given by Lemma 7.6, which depends on Assumptions 3.4 and 3.5, is enough for the argument above to follow as Equation 7.12 holds with the limiting transition rates $q^{n}_{\xi\eta}$ replaced by the finite $N$ transition rates $q^{N,n}_{\xi\eta}$, which follows by Assumption 3.7 (here $\xi,\eta$ are simply placeholders for the true elements used in the above equation). (7.8) implies the convergence of $\bar{G}^{N}$ in $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$ to $G_{\tilde{Q},\lambda}^{n}$ by (EK, 09, Corollary 8.9, Chapter 4) and (7.6).
The discrete-time ancestral graph $G^{N,n}=(G^{N,n}(k))_{k\in\mathbb{Z}_{+}}$ can be viewed graphically as having $n$ initial nodes (the nodes at the bottom of Figure 7). These $n$ nodes correspond to the $n$ individuals $\{\hat{X}_{i}^{N}(0)\}_{1\leq i\leq n}$ from whom we sampled the $n$ lineages at time-step 0. Viewed as a $\Omega$-valued process, this graph has the following key advantages: it get rid of the labels of all individuals, while keeping track of the set of all possible ancestral individuals (of the sample) and their parental relationships.
@@
Define $x_{i}^{N}(0):=i$ for all $1\leq i\leq n$. Suppose $x_{i}^{N}(j)$ is defined for all $j\leq k$ for induction. If $\hat{X}_{i}^{N}(k)$ experiences an outcrossing event in the $k$th time-step, then $x_{i}^{N}(k+1)$ is equal to $l^{N}(k+1)$ with probability $\frac{1}{2}$ or else remains equal to $x_{i}^{N}(k)$ with probability $\frac{1}{2}$. (In the event that there is more than one outcrossing event in a single time-step, it may be that each outcrossing individual coalesces in the position $l+1$; in the limited outcrossing regime, this possibility is negligible, so we do not need to consider this case in any detail.) We therefore have $x_{i}^{N}=\left(x_{i}^{N}(k)\right)_{k\in\mathbb{Z}_{+}}$ defined for all time. If any two walks $x_{i}^{N}$ and $x_{j}^{N}$ coalesce at any point in time, they are then required to remain together for all time thereafter. This gives a family of $n$ simple random walks on $G^{N,n}$ satisfying
@@
We may also define an associated coalescent process based on these $x_{i}^{N}$.
@@
Let $\Pi_{G}^{N,n}=\left(\Pi_{G}^{N,n}(t)\right)_{\in\mathbb{R}_{+}}$ be the sped-up coalescent process defined by
$i\sim j\text{ with respect to }\Pi_{G}^{N,n}(t)\text{ if }x_{i}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right)=x_{j}^{N}\left(\lfloor tc_{N}^{-1}\rfloor\right),$
and define the quenched law
@@
Note that $\mathcal{L}^{N,n}_{G}$ defined in (7.19) is equal to $\mathbb{P}\left(\chi^{N,n}(\lfloor tc_{N}^{-1}\rfloor)\in\cdot\,\mid\,\mathcal{A}_{N}\right)$ which admits no random time change, unlike that of $\mathcal{L}^{N,n}$. This is because, unlike the random walks on the pedigree, we assume being in the same node in the ancestral graph $G^{N,n}$ means instantaneous coalescence of the particles. We will show in Lemma 7.14 that $\mathcal{L}^{N,n}$ and $\mathcal{L}^{N,n}_{G}$ have the same weak limit as $N\to\infty$.
A suitable notion of continuity for random walks on $\Omega$-valued processes, established in Section 7.3, will then give the claimed result Theorem 4.7.
7.3 Coalescing random walks on $\Omega$-valued processes
To each element $g=\left(g(t)\right)_{t\in\mathbb{R}_{+}}=\left(l(t),m(t),\xi(t)\right)_{t\in\mathbb{R}_{+}}$ in the Skorokhod space $\mathcal{D}\left(\mathbb{R}_{+},\Omega\right)$, we will define an associated coalescent process $\chi(g)$. Let $g(t-)=\lim_{s\uparrow t}g(s)=(l(t-),m(t-),\xi(t-))$ denote the left limit of $g$ at $t$. We suppose that $g(0)=(l_{0},m_{0},\xi_{0})$ and define, for all $1\leq i\leq l_{0}$ that $x_{i}(0)=i$. The $x_{i}$ may be extended to all time as follows: